OpenStudy (anonymous):

What is the equation for the ellipse with horizontal minor axis of length 8 and foci located at (0, 5) and (0, −5).

OpenStudy (amistre64):

the center is (0,0); along the y axis the minor spread is 4 x^2 y^2 ---- + ---- = 1 16 a^2 a = the hyp of a right triangle with legs b and foci distance from center (5) a^2 = b^2 + c^2 a^2 = 16 + 25 = 41 x^2 y^2 ---- + ---- = 1 ; it shoud be lol 16 41